OpenStudy (anonymous):

If A is the arithmetic mean between two numbers and G is geometric mean, find the two numbers.

OpenStudy (anonymous):

Do u knw hw to find AM and GM

OpenStudy (anonymous):

a+b/2 and sqt(ab)

OpenStudy (anonymous):

and AM > = GM

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a + b = 2A ab = G^2 Find a and b ...) r\

OpenStudy (anonymous):

a + G^2/a = 2A ?

OpenStudy (anonymous):

ok gotcha ty ! :D